# Almost Surely Asymptotically Constant Graph Neural Networks

### v2

@@ 6.1 Empirical results on random graph models
One striking feature of Figure 2 is that each given model converges to the same outputs on both the dense and the logarithmic growth distributions, while on the sparse distributions the eventual constant output varies (Q3). We can relate this to the distinct proof strategy employed in the sparse case, which uses convergence of the proportions of local isomorphism types. We can have many local isomorphism types, and the experiments show that what we converge to depends on the proportion of these. In the other cases, regardless of the distribution, the neighbourhood sizes are unbounded, so there is asymptotically almost surely one ‘local graph type’.
@@ 7 Discussion and limitations
We have shown there is a wide convergence phenomenon for probabilistic classifiers expressed even in very advanced GNN architectures, and it applies to a great variety of random graph models. Rather than having separate proof techniques per GNN model, our paper introduces a broad language where such models can be situated, and provides techniques at the level of term languages.
We emphasize that a.a.s. convergence is not universal, and it does not hold even for ER with arbitrary $p(n)$ going to 00. It is easy to see that if we alternate between the sparse case and the log or root case, we will get divergence. A.a.s. convergence also depends strongly on the details of the term language: if we add sum aggregation, the term values clearly diverge, and similarly if we allow non-smooth functions – e.g. testing whether a term is strictly above a constant. A.a.s. convergence does not hold if max aggregation is added, although we believe that weaker convergence laws can be established for this aggregator.
@@ C.1 Combinatorial and growth lemmas
At various points in the following we make statements concerning particular nodes of an Erdős-Rényi graph which hold across all such graphs. For example, we state that for any $u$ we have that its degree $d(u)$ is the sum of $n-1$ Bernoulli random variables. To make sense of such statements we can fix a canonical enumeration of the nodes of every graph and view our node meta-variables as ranging over indices. So we would translate the previous example as “for any node index $u$ the degree of the $u$th node in $\mathrm{ER}(n,p(n))$ is the sum of $n-1$ Bernoulli random variables.”
@@
Recall the Log Neighbourhoods Lemma, Lemma 5.3. We prove a stronger result, which allows us finer control over the growth in each of the non-sparse cases.
@@
Take any $\wedge$-desc $\phi$ on $k$ variables, and choose $i\leq k$. We consider $k$-tuples of nodes $\bar{u}$ satisfying $\phi$ and the degrees of the $i$th node $u_{i}$.
@@
    Now, any $d(u)$ is the sum of $n-1$ i.i.d. Bernoulli random variables with parameter $p(n)$. Hence by Hoeffding’s Inequality (Theorem B.4) and a union bound, there is $N_{2}$ such that for all $n\geq N_{2}$, with probability at least $1-\theta$, for every node $u$ we have:
    $\left|\frac{d(u)}{n}-\tilde{p}\right|\geq\epsilon$
@@
The basic form of the following “regularity lemma” states that for every $k\in\mathbb{N}^{+}$ and $i\leq k$, whenever we have a large subset $S$ of the $(k+1)$-tuples $(\bar{u},v)$ with each $u_{i}$ adjacent to $v$, there is a large subset of the $k$-tuples $\bar{u}$ such that each $u_{i}$ is adjacent to a large set of $v$’s such that $(\bar{u},v)\in S$. The full regularity lemma states that this is the case also when we restrict the tuples $\bar{u}$ to a certain $\wedge$-desc.
@@
Let $p\colon\mathbb{N}\to[0,1]$ satisfy one of the conditions other than sparsity. Take a $\wedge$-desc $\phi$ on $k$ variables and fix $i\leq k$. Then for every $\gamma,\theta>0$ there is $N\in\mathbb{N}$ and $\delta^{\prime}>0$ such that for all $n\geq N$ and $\delta<\delta^{\prime}$, with probability at least $1-\theta$ we have that, whenever $S\subseteq\{(\bar{u},v)\mid\phi(\bar{u})\wedge u_{i}Ev\}$ is such that:
@@
Let $\mathbb{F}$ be the probability distribution of the node features, which, for the sake of convenience, we assume to have domain $[0,1]^{d}$. Note that in each case the probability function $p(n)$ converges to some $\tilde{p}$ (which is 00 in the root growth and logarithmic growth cases).
Recall from the body of the paper that a $(\bar{u},d)$ feature-type controller is a Lipschitz function taking as input pairs consisting of a $d$-dimensional real vector and a $\bar{u}$ graph type. Recall from Lemma 5.4 that for each term $\pi$, we need to define a controller $e_{\pi}$ that approximates it. We first give the construction of $e_{\pi}$ and then recall and verify what it means for $e_{\pi}$ to approximate $\pi$.
@@
When $\pi={\bm{c}}$ a constant define:
@@
Note that in both the root growth and logarithmic growth cases, $\alpha(t,t^{\prime})$ is non-zero precisely when $t^{\prime}$ specifies no edges between $\bar{u}$ and $v$.
@@
Consider $\pi=\sum_{v\in u_{i}}\rho(\bar{u},v)\star h(\eta(\bar{u},v))$. Define:
@@
With the controllers $e_{\pi}$ defined, we now prove that they satisfy the inductive lemma (Lemma 5.4). We restate this here using our slightly modified notation.
@@
For every $\epsilon,\delta,\theta>0$ and $\wedge$-desc $\psi(\bar{u})$, there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ in the space of graphs of size $n$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\delta$ satisfy:
@@
    Take $\epsilon,\delta,\theta>0$ and $\wedge$-desc $\psi(\bar{u})$. By the induction hypothesis there is $N\in\mathbb{N}$ such that for all $n\geq N$ and for every $i\leq r$, with probability at least $1-\theta$, we have that out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\delta$ satisfy:
@@
    $f(e_{\rho_{1}}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u})),\dots,e_{\rho_{r}}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u})))$
@@
    is at most $L_{f}\epsilon$, where $L_{f}$ is the Lipschitz constant of $f$.
@@
    Take $\epsilon,\delta,\theta>0$. Take a $\wedge$-desc $\psi(\bar{u})$.
@@
    By Lemma C.4 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\gamma$ satisfy that for at least $1-(\delta+\delta^{2})$ of the $v$’s satisfy equation ($\dagger_{\rho}$) and equation ($\dagger_{\eta}$).
@@
    Let $B_{n}$ be the proportion out of all tuples $\bar{u}$ for which $\psi(\bar{u})$ holds such that:
@@
    Then by linearity of expectation $\operatorname*{\mathbb{E}}[B_{n}]\leq\theta\delta$ for all $n\geq N_{4}$, and hence by Markov’s Inequality (LABEL:thm:Markov'sInequality):
@@
    We proceed similarly to the global weighted mean case, this time making use of the conditioning $\wedge$-desc in the inductive invariant. Indeed, notice that when we apply the inductive invariant below, we add $u_{i}Ev$ to our condition.
    Take $\epsilon,\delta,\theta>0$. Take a $\wedge$-desc $\psi(\bar{u})$.
@@
    By Lemma C.4 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})\wedge\exists v\colon u_{i}Ev$, at least $1-\gamma$ satisfy that for at least $1-(\delta+\delta^{2})$ of the $v$ for which $\psi(\bar{u})\wedge u_{i}Ev$ holds also satisfy equation ($\dagger_{\rho}^{\textrm{l}oc}$) and equation ($\dagger_{\eta}^{\textrm{l}oc}$).
@@
    is at least $1-\delta$. Then, as before, there is $N_{4}\geq N_{3}$ such that for all $n\geq N_{4}$, probability at least $1-\theta$, for all pairs $(t,t^{\prime})$ the proportion of all tuples $\bar{u}$ for which $\psi(\bar{u})$ and $t(\bar{u})$ hold such that:
@@
    Using a similar argument to the global weighted mean case, there is $N_{5}$ such that for all $n\geq N_{5}$, with probability at least $1-\theta$ for all $t^{\prime}$ such that $\alpha(t,t^{\prime})>0$ we have that for every at least $1-\delta$ of the tuples $\bar{u}$ such that $\psi(\bar{u})$ and $t(\bar{u})$ hold:
@@
    Proceeding as before then, for all $n\geq\max(N_{3},N_{4},N_{5})$, with probability at least $1-\theta$ for at least $1-\delta$ of the tuples $\bar{u}$ for which $\psi(\bar{u})$ holds we have that:
@@
    This completes the proof of Lemma 5.4.∎
@@ D.1 Combinatorial tools for the sparse case
Consider sampling a graph $G$ from either the sparse Erdős-Rényi or Barabási-Albert distributions. Let $(T,(\bar{w},y))$ be a $(k+1)$-rooted graph and $\ell\in\mathbb{N}$. There is $q_{T}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all $k$-tuples of nodes $\bar{u}$:
||$\left|\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,(\bar{w},y))\}|}{n}-q_{T}\right|<\epsilon$
This follows from two important properties shared by these distributions. The first is what’s known as “weak local convergence”, which is essentially a non-parametrised version of Theorem D.3.
@@
See [51, Theorem 2.18] for the sparse Erdős-Rényi distribution and [17, Theorem 4.2.1] for the Barabási-Albert distribution. ∎
The second property is that the maximum neighbourhood sizes of nodes grows slower than $n$.
Lemma D.6.
Consider sampling a graph $G$ from either the sparse Erdős-Rényi or Barabási-Albert distributions. Then for every $\ell\in\mathbb{N}$ and $\gamma>0$ there is $N>N_{2}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all nodes $u$:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(u)|}{n}<\gamma$
Proof.
For the Barabási-Albert distribution we appeal to a well-known result that if $M_{n}$ is the maximum degree of the $n$-vertex Barabási-Albert graph then for every $\epsilon>0$ we have that:
$\mathbb{P}(M_{n}<n^{\frac{1}{2}+\epsilon})\to 1\text{ as }n\to\infty$
See [41, 38, 39].
Now consider the sparse Erdős-Rényi case. Let $p(n)=Kn^{-1}$. Take any node $u$. Then $d(u)$ is the sum of $n-1$ independent Bernoulli random variables with parameter $p(n)$. Hence by Chernoff’s Inequality (Theorem B.3):
$\mathbb{P}\left(\left|d(u)-p(n)\right|>\log(n)\right)$ $<e^{-K}\left(\frac{eK}{\log(n)}\right)^{\log(n)}$
$\quad=e^{-K}n^{1+\log(K)-\log(\log(n))}$
There is $N_{1}$ such that for all $n\geq N_{1}$ we have:
$21+\log(K)-\log(\log(n))<-$ 2
Hence, taking a union bound, there is $N_{2}$ such that for all $n\geq N_{2}$, with probability at least $1-\theta$ we have that for all nodes $u$:
$\left|d(u)-p(n)\right|<\log(n)$
Condition on the event that this happens. Then for every $u$ we have:
$|\operatorname{\mathcal{N}}^{G}_{\ell}(u)|<\log(n)^{\ell}$
So:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(u)|}{n}<\frac{\log(n)^{\ell}}{n}$
and the result follows. ∎
With these two properties we can prove our combinatorial result.
@@
There are two cases.
@@
There is a path from $y$ to some node $w_{i}$ of $\bar{w}$ in $T$.
Set $q_{T}=0$. Note that in this case, for all tuples $\bar{u}$:
$\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}\subseteq\operatorname{\mathcal{N}}^{G}_{\ell}(u_{i})$
By Lemma D.6, for every $\epsilon>0$ there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all nodes $v$:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(v)|}{n}<\epsilon$
Hence, for all tuples $\bar{u}$:
||$\left|\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}|}{n}\right|<\left|\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(u_{i})|}{n}\right|<\epsilon$
This completes the proof for Case 1.
@@
There is no path from $y$ to any node of $\bar{w}$ in $T$.
Then for all tuples $\bar{u}$:
$\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}=\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}\setminus\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})$
@@
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}|}{n}-\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})|}{n}\leq\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}|}{n}$

### v3

@@ 6.1 Empirical results on random graph models
One striking feature of Figure 3 is that the eventual constant output of each model is the same for the dense and logarithmic growth distributions, but not for the sparse distribution (Q3). We can relate this to the distinct proof strategy employed in the sparse case, which uses convergence of the proportions of local isomorphism types. There are many local isomorphism types, and the experiments show that what we converge to depends the proportion of these. In all other cases the neighbourhood sizes are unbounded, so there is asymptotically almost surely one ‘local graph type’.
@@ 7 Discussion and limitations
We have demonstrated a wide convergence phenomenon for real-valued classifiers expressed even in very advanced GNN architectures, and it applies to a great variety of random graph models. Rather than having separate proof techniques per GNN model, our paper introduces a broad language where such models can be situated, and provides techniques at the level of term languages. Although our top-level theorems deal with graph-level tasks, along the way we provide strong limitative results on what can be achieved on random graphs for node- or edge-level real-valued tasks: see Theorem 5.3.
The principal limitations of our work come our assumptions. In particular, we assume that the initial node embeddings are i.i.d. This assumption is used in the application concentration inequalities throughout the proofs, so loosening it would require careful consideration.
Our main results show that many GNN architectures cannot distinguish large graphs. To overcome this limitation, one could consider moving beyond our term language. For example, if we add sum aggregation, the term values clearly diverge, and similarly if we allow non-smooth functions, such as linear inequalities. Further we emphasize that a.a.s. convergence is not universal for our term language, and it does not hold even for ER with arbitrary $p(n)$ going to 00: see Theorem 5.7.
@@ C.1 Combinatorial and growth lemmas
At various points in the following we will make statements concerning particular nodes of an Erdős-Rényi graph without fixing a graph. For example, we state that for any $u$ we have that its degree $d(u)$ is the sum of $n-1$ Bernoulli random variables. To make sense of such statements we can fix a canonical enumeration of the nodes of every graph and view our node meta-variables as ranging over indices. So we would translate the previous statement as “for any node index $u$ the degree of the $u$th node in $\mathrm{ER}(n,p(n))$ is the sum of $n-1$ Bernoulli random variables.”
Both the combinatorial and growth lemmas, and the main inductive results themselves require saying that certain events hold with high probability. Inductively, we will need strengthenings of these compared with the statements given in the body. We will need to consider conditional probabilities, where the conditioning is on a conjunction of atomic statements about the nodes.
Definition C.3.
A $\wedge$-description $\phi(\bar{u})$ is a conjunction of statements about the variables $\bar{u}$ of the form $u_{i}Eu_{j}$, which expresses that there exists an edge between $u_{i}$ and $u_{j}$. A $\wedge$-description $u_{i_{1}}Eu_{j_{1}}\wedge\cdots\wedge u_{i_{m}}Eu_{j_{m}}$ holds if and only if there is an edge between $u_{i_{\ell}}$ and $u_{j_{\ell}}$ for each $\ell$.
The reason for strengthening the results in this way will become clearer when we are proving the local weighted mean induction step $\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}\rho(\bar{u},v)\star h(\eta(\bar{u},v))$. There we will need our auxiliary results to hold conditioned on $vEu_{i}$, along with any additional requirements coming from previous induction steps.
@@
We will need to know that in the non-sparse case, neighborhoods are fairly big:
Lemma C.4 (Log neighbourhood bound).
We prove a stronger result, which allows us finer control over the growth in each of the non-sparse cases. The strengthening will involve the $\wedge$-descriptions of Definition C.3.
@@
Take any $\wedge$-description $\phi$ on $k$ variables, and choose $i\leq k$. We consider $k$-tuples of nodes $\bar{u}$ satisfying $\phi$ and the degrees of the $i$th node $u_{i}$.
@@
    Now, any $d(u)$ is the sum of $n-1$ i.i.d. Bernoulli random variables with parameter $p(n)$: see Remark C.2 for the formalization. Hence by Hoeffding’s Inequality (Theorem B.4) and a union bound, there is $N_{2}$ such that for all $n\geq N_{2}$, with probability at least $1-\theta$, for every node $u$ we have:
    $\left|\frac{d(u)}{n}-\tilde{p}\right|<\epsilon$
@@
The basic form of the following “regularity lemma” states that for every $k\in\mathbb{N}^{+}$ and $i\leq k$, whenever we have a large subset $S$ of the $(k+1)$-tuples $(\bar{u},v)$ with each $u_{i}$ adjacent to $v$, there is a large subset of the $k$-tuples $\bar{u}$ such that each $u_{i}$ is adjacent to a large set of $v$’s such that $(\bar{u},v)\in S$. The full regularity lemma states that this is the case also when we restrict the tuples $\bar{u}$ to a certain $\wedge$-description.
@@
Let $p\colon\mathbb{N}\to[0,1]$ satisfy one of the conditions other than sparsity. Take a $\wedge$-description $\phi$ on $k$ variables and fix $i\leq k$. Then for every $\gamma,\theta>0$ there is $N\in\mathbb{N}$ and $\delta^{\prime}>0$ such that for all $n\geq N$ and $\delta<\delta^{\prime}$, with probability at least $1-\theta$ we have that, whenever $S\subseteq\{(\bar{u},v)\mid\phi(\bar{u})\wedge u_{i}Ev\}$ is such that:
@@
Let $\mathbb{F}$ be the probability distribution of the node features, which, for the sake of convenience, we assume to have domain $[0,1]^{d}$. The domain $[0,1]^{d}$ can be replaced by the more general domain $[a,b]^{d}$ in the results and proofs without further modification. But we note that the domain $[0,1]^{d}$ is already sufficient for the results to hold in the general case. Indeed, suppose $\tau(\bar{x})$ is a term which we apply to a graph distribution $\mathcal{D}$ with features in $[a,b]^{d}$. Modify this distribution to $\mathcal{D}^{\prime}$ by applying the function $\bar{z}\mapsto(\bar{z}-a)/(b-a)$ to the features. This is now a distribution with features in $[0,1]^{d}$. Modify $\tau$ to $\tau^{\prime}$ by replacing each $\mathrm{H}(x)$ by $F(\mathrm{H}(x))$, where $F$ is the function $\bar{z}\mapsto(b-a)\bar{z}+a$. Then evaluating $\tau^{\prime}$ on $\mathcal{D}^{\prime}$ is equivalent to evaluating $\tau$ on $\mathcal{D}$.
Note that in each case the probability function $p(n)$ converges to some $\tilde{p}$ (which is 00 in the root growth and logarithmic growth cases).
Recall from the body of the paper that a $(\bar{u},d)$ feature-type controller is a Lipschitz function taking as input pairs consisting of a $d$-dimensional real vector and a $\bar{u}$ graph type. Recall from Theorem 5.3 that for each term $\pi$, we need to define a controller $e_{\pi}$ that approximates it. We first give the construction of $e_{\pi}$ and then recall and verify what it means for $e_{\pi}$ to approximate $\pi$.
@@
When $\pi={\bm{c}}$, a constant, define:
@@
Note that in both the root growth and logarithmic growth cases, $\alpha(t,t^{\prime})$ is non-zero precisely when $t^{\prime}$ specifies no edges between $\bar{u}$ and $v$. These relative atomic type weights will play a key role in the construction.
@@
Note that when $\tilde{p}=0$ (as in the root growth and logarithmic growth cases), we have that:
$\alpha(t,t^{\prime})=\begin{cases}1&\text{if $t^{\prime}$ specifies no edges between $\bar{u}$ and $v$}\\ 0&\text{otherwise}\end{cases}$
Therefore the controller becomes, letting $t^{\varnothing}$ be the type with no edges between $\bar{u}$ and $v$:
$e_{\pi}(\bar{\bm{a}},t)=\frac{\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(\bar{\bm{a}},{\bm{b}},t^{\varnothing})h(e_{\eta}(\bar{\bm{a}},{\bm{b}},t^{\varnothing}))\right]}{\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[h(e_{\eta}(\bar{\bm{a}},{\bm{b}},t^{\varnothing}))\right]}$
@@
Consider $\pi=\sum_{v\in N(u_{i})}\rho(\bar{u},v)\star h(\eta(\bar{u},v))$. Define:
@@
With the controllers $e_{\pi}$ defined, we now prove that they satisfy Theorem (Theorem 5.3). We state the strengthened version of this result here using the notation of $\wedge$-decriptions from Definition C.3:
@@
For every $\epsilon,\delta,\theta>0$ and $\wedge$-description $\psi(\bar{u})$, there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ in the space of graphs of size $n$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\delta$ satisfy:
@@
    Take $\epsilon,\delta,\theta>0$ and $\wedge$-description $\psi(\bar{u})$. By the induction hypothesis there is $N\in\mathbb{N}$ such that for all $n\geq N$ and for every $i\leq r$, with probability at least $1-\theta$, we have that out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\delta$ satisfy:
@@
    $e_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))=f(e_{\rho_{1}}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u})),\dots,e_{\rho_{r}}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u})))$
@@
    is at most $L_{f}\epsilon$, where $L_{f}$ is the Lipschitz constant of $f$. Since $L_{f}$ is a constant, this case follows.
@@
    Take $\epsilon,\delta,\theta>0$. Take a $\wedge$-description $\psi(\bar{u})$.
@@
    By Lemma C.6 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})$, at least $1-\gamma$, at least $1-(\delta+\delta^{2})$ of the $v$’s satisfy equation ($\dagger_{\rho}$) and equation ($\dagger_{\eta}$).
@@
    $\left\lVert\frac{1}{n}\sum_{v\mid t^{\prime}(\bar{u},v)}f^{\circ}_{\pi}(\bar{u},v)-\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{n}\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(H_{G}(\bar{u},{\bm{b}}),t^{\prime})h(e_{\eta}(H_{G}(\bar{u},{\bm{b}}),t^{\prime}))\right]\right\rVert\geq\frac{1}{2^{k+1}}\epsilon$ ($\heartsuit$)
@@
    $2d\exp\left(-\frac{\epsilon^{2}|\llbracket t^{\prime}(\bar{u})\rrbracket|}{2\Lambda^{2}}\right)<\theta\delta$ ($\clubsuit$)
    Let $B_{n}$ be the proportion, out of all tuples $\bar{u}$ for which $\psi(\bar{u})$ holds [^3] Recall that $\psi$ is the $\wedge$-description on which we condition the tuples $\bar{u}$ in the inductive invariant., such that:
@@
[…]
    Then by linearity of expectation, $\operatorname*{\mathbb{E}}[B_{n}]\leq\theta\delta$ for all $n\geq N_{4}$, and hence by Markov’s Inequality (LABEL:thm:Markov'sInequality):
@@
    We proceed similarly to the global weighted mean case, this time making use of the conditioning $\wedge$-description in the inductive invariant. Indeed, notice that when we apply the inductive invariant below, we add $u_{i}Ev$ to our condition.
    Take $\epsilon,\delta,\theta>0$. Take a $\wedge$-description $\psi(\bar{u})$.
@@
    By Lemma C.6 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$ the following event happens. Out of all the tuples $\bar{u}$ such that $\psi(\bar{u})\wedge\exists v\colon u_{i}Ev$, a proportion of at least $1-\gamma$ of them satisfy the following. At least $1-(\delta+\delta^{2})$ of the nodes $v$ for which $\psi(\bar{u})\wedge u_{i}Ev$ holds also satisfy equation ($\dagger_{\rho}^{\textrm{l}oc}$) and equation ($\dagger_{\eta}^{\textrm{l}oc}$).
@@
    is at least $1-\delta$. Then, as before, there is $N_{4}\geq N_{3}$ such that for all $n\geq N_{4}$, with probability at least $1-\theta$, for all pairs $(t,t^{\prime})$ the proportion of all tuples $\bar{u}$ for which $\psi(\bar{u})$ and $t(\bar{u})$ hold such that:
@@
    Using a similar argument to the global weighted mean case, there is $N_{5}$ such that for all $n\geq N_{5}$, with probability at least $1-\theta$ for all $t^{\prime}$ such that $\alpha(t,t^{\prime})>0$ we have that for at least $1-\delta$ of the tuples $\bar{u}$ such that $\psi(\bar{u})$ and $t(\bar{u})$ hold:
@@
    We now proceed as in the global weighted mean case, creating a random variable similar to $B_{n}$ and making using of linearity of expectation. We get that for all $n\geq\max(N_{3},N_{4},N_{5})$, with probability at least $1-\theta$ for at least $1-\delta$ of the tuples $\bar{u}$ for which $\psi(\bar{u})$ holds we have that:
@@
    This completes the proof of Theorem 5.3.∎
@@ D.1 Combinatorial tools for the sparse case
Definition D.3.
Let $(T,\bar{w})$ be a $k$-rooted graph and let $(T^{\prime},\bar{w},y)$ be a $(k+1)$-rooted graph. Then $T^{\prime}$ extends $T$ if $T$ is a subgraph of $T^{\prime}$. Let $\mathsf{Ext}(T)$ be the set of all $(k+1)$-rooted graphs extending $(T,\bar{w})$.
@@
Consider sampling a graph $G$ from either the sparse Erdős-Rényi or Barabási-Albert distributions. Let $(T,\bar{w})$ be a $k$-rooted graph and take $\ell\in\mathbb{N}$. Then for every $(k+1)$-rooted graph $(T^{\prime},\bar{w},y)$ which extends $(T,\bar{w})$ there is $q_{T^{\prime}\mid T}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all $k$-tuples of nodes $\bar{u}$ such that $\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})\cong(T,\bar{w})$:
||$\left|\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}|}{n}-q_{T}\right|<\epsilon$
Moreover, we have that:
$\sum_{T^{\prime}\in\mathsf{Ext}(T)}q_{T^{\prime}\mid T}=1$
We refer to the $q_{T^{\prime}\mid T}$ as relative neighborhood weights. They will play a role in defining the controllers, analogous to that of the relative atomic type weights $\alpha(t,t^{\prime})$ in the non-sparse case.
Theorem D.4 follows from what is known as “weak local convergence” in the literature. It is essentially a non-parametrised version of Theorem D.4.
@@
See [48, Theorem 2.18] for the sparse Erdős-Rényi distribution and [17, Theorem 4.2.1] for the Barabási-Albert distribution. ∎
@@
Take $T^{\prime}\in\mathsf{Ext}(T)$. There are two cases.
@@
There is a path from $y$ to some node $w_{i}$ of $\bar{w}$ in $T^{\prime}$.
Set $q_{T^{\prime}\mid T}=0$. Note that in this case, for all tuples $\bar{u}$ such that $\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})\cong T$ we have that:
$\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}\subseteq\operatorname{\mathcal{N}}^{G}_{\ell}(u_{i})\}$
Since $\operatorname{\mathcal{N}}^{G}_{\ell}(u_{i})$ is determined by $T$ and is thus of fixed size, the proportion:
||$\left|\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}|}{n}\right|$
tends to 00 as $n$ grows. This completes the proof for Case 1.
@@
There is no path from $y$ to any node of $\bar{w}$ in $T^{\prime}$. Then for all tuples $\bar{u}$ such that $\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})\cong T$ we have that:
$\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}=\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)\}\setminus\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})$
@@
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)\}|}{n}-\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})|}{n}\leq\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}|}{n}$